# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v1

@@ 3.4 Putting it together
We can now show that $\mathsf{In\text{-}and\text{-}Out}$ subsumes previous results on uniform sampling from convex bodies (such as $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$), providing detailed versions of the main results in Section 1.2.
We first establish that the query complexity of $\mathsf{In\text{-}and\text{-}Out}$ matches that of the $\mathsf{Ball\ walk}$ under stronger divergences. Recall that $2\|\cdot\|_{\mathsf{TV}}^{2}\leq\mathsf{KL}\leq\log(1+\chi^{2})\leq\chi^{2}$.
@@
  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}\log\frac{M}{\eta\varepsilon})$ iterations. With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{\mathcal{O}}(qMd^{2}\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}(\log\frac{1}{\eta\varepsilon})^{5})$ expected number of membership queries in total.
  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{\mathcal{O}}(qd^{2}\log\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{\mathcal{O}}(qMd^{2}(\log\frac{1}{\eta\varepsilon})^{5})$ membership queries in expectation.
@@
We just put together Lemma 14 and Theorem 23. For target accuracy $\varepsilon>0$, we use the $\mathcal{R}_{q}$-decay under (7) for $q\geq 2$ in Theorem 23. The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$. Due to $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}\,\log d)$ (Lemma 18), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}\log\frac{M}{\eta\varepsilon})$ iterations. Since each iteration has $\eta/m$-failure probability by Lemma 14, the union bound ensures that the total failure probability is at most $\eta$ throughout $m$ iterations. Lastly, each iteration requires $\widetilde{\mathcal{O}}(M(\log\frac{1}{\eta\varepsilon})^{4})$ membership queries in expectation by Lemma 14. Therefore, $\mathsf{In\text{-}and\text{-}Out}$ uses $\widetilde{\mathcal{O}}(qMd^{2}\min(D^{2},\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}})(\log\frac{1}{\eta\varepsilon})^{5})$ expected number of membership queries over $m$ iterations. Since $\mathcal{R}_{q}$ is non-decreasing in $q$, we can obtain the desired bound on $\mathcal{R}_{q}$ for $q\in[1,2)$.
For isotropic $\pi^{X}$, we have $\mathrm{Cov}(\pi^{X})=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\log d)$ (see Lemma 18). ∎
@@
$\mu_{1}^{X}(x)=\mathds{1}_{\mathcal{K}}(x)\cdot\int\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\exp\bigl{(}-\frac{1}{2h}\|y-x_{0}\|^{2}\bigr{)}}{(2\pi h)^{d/2}\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\,\mathrm{d}x}\,\mathrm{d}y\,.$
By Young’s inequality, $\|y-x\|^{2}\leq(\|y\|+D)^{2}\leq\frac{3}{2}\|y\|^{2}+3D^{2}$ for $x\in\mathcal{K}$. Hence,
$\int\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\exp\bigl{(}-\frac{1}{2h}\|y-x_{0}\|^{2}\bigr{)}}{\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\,\mathrm{d}x}\,\mathrm{d}y$
$\leq$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\bigl{(}\|y-x\|^{2}+\|y-x_{0}\|^{2}-\frac{3}{2}\|y\|^{2}\bigr{)}\Bigr{)}\,\mathrm{d}y$
$=$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\bigl{(}\frac{1}{2}\|y-2(x+x_{0})\|^{2}+(\|x\|^{2}+\|x_{0}\|^{2}-2\|x+x_{0}\|^{2})\bigr{)}\Bigr{)}\,\mathrm{d}y$
$\leq$ $\frac{\exp(5h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{4h}\|y-2(x+x_{0})\|^{2}\Bigr{)}\,\mathrm{d}y$
@@
For $h=\tilde{\Theta}(d^{-2})$, $\mathsf{In\text{-}and\text{-}Out}$ requires $\widetilde{\mathcal{O}}(qd^{2}D^{2})$ iterations and in particular $\widetilde{\mathcal{O}}(qd^{2}D)$ iteration for isotropic uniform distributions. These results match those of $\mathsf{Speedy\ walk}$ [kannan2006blocking, lee2017eldan] (see Theorem 31).
@@ Acknowledgements.
We are deeply grateful to Andre Wibisono and Sinho Chewi for helpful comments and pointers to the literature for Lemma 19. This work was supported in part by NSF award 210644, NSERC through the CGS-D award, and a Simons Investigator award.
[…]
@@ Appendix A Related work
Sampling from constrained log-concave distributions is a fundamental task arising in many fields. Uniform sampling with convex constraints is its simplest manifestation, which was first studied as a core subroutine for a randomized volume-computation algorithm [dyer1991random]. Since then, this fundamental problem has been studied for over three decades [lovasz1990mixing, lovasz1993random, kannan1997random, lovasz2003hit, lovasz2006hit, bubeck2018sampling, pmlr-v65-brosse17a]. We review these algorithms, grouping them under three categories — geometric random walks, structured samplers, and diffusion-type samplers. Below, $\mathcal{K}$ is convex.
@@
We discuss two geometric random walks – $\mathsf{Ball\ walk}$ [lovasz1993random, kannan1997random] and $\mathsf{Hit\text{-}and\text{-}Run}$ [smith1984efficient, lovasz1999hit].
$\mathsf{Ball\ walk}$ is a simple metropolized random walk; it draws $y$ uniformly at random from a ball of radius $\delta$ centered at a current point $x$, and moves to $y$ if $y\in\mathcal{K}$ and stays at $x$ otherwise. In the literature, $\mathsf{Ball\ walk}$ actually refers to a composite algorithm consisting of [$\mathsf{Speedy\ walk}+$ rejection sampling], where $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$ (see Appendix B for details). The step size $\delta$ should be set to $\mathcal{O}(d^{-1/2})$ to avoid stepping outside of $\mathcal{K}$. [kannan1997random] showed that $\mathsf{Ball\ walk}$ needs $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ membership queries to be $\varepsilon$-close to $\pi^{\mathcal{K}}$ in $\mathsf{TV}$, where $D$ is the diameter of $\mathcal{K}$, and the warmness parameter $M$ measures the closeness of the initial distribution to the target uniform distribution $\pi^{\mathcal{K}}$.
$\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$. [lovasz2006hit] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{X\sim\pi^{\mathcal{K}}}[\|X-\mathbb{E}[X]\|^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $\mathcal{O}(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries. Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to $\mathsf{Ball\ walk}$.
Both algorithm are affected by skewed shape of $\mathcal{K}$ (i.e., large $D$ or $R$), so these samplers are combined with pre-processing step called rounding. This procedure finds a linear transformation that makes the geometry of $\mathcal{K}$ less skewed and so more amenable to sampling. In literature, there exists a randomized algorithm [jia2021reducing] that rounds $\mathcal{K}$ and generates a good warm start (i.e., $M=\mathcal{O}(1)$), with $\mathsf{Ball\ walk}$ used as a core subroutine. This algorithm takes up $\widetilde{\mathcal{O}}(d^{3})$ queries in total, and in such position with the good warm start, $\mathsf{Ball\ walk}$ only needs $\widetilde{\mathcal{O}}(d^{2}\log\frac{1}{\varepsilon})$ queries to sample from $\pi^{\mathcal{K}}$.
@@
The first canonical example of such a zeroth-order sampler is $\mathsf{Dikin\ walk}$ used when $\mathcal{K}$ is given by $m$ linear constraints [kannan2012random]; it draws a uniform sample from an ellipsoid (characterized by $\nabla^{2}\phi$) of fixed radius around a current point, and is often combined with a Metropolis adjustment. [kannan2012random] shows that $\mathsf{Dikin\ walk}$ mixes in $\mathcal{O}(md\log\frac{M}{\varepsilon})$ steps, although each iteration is slightly more expensive than one membership query. This algorithm requires no rounding, but still needs a good warm-start, which can be achieved by an annealing-type algorithm using $\widetilde{\mathcal{O}}(md)$ iterations of $\mathsf{Dikin\ walk}$ [kook2024gaussian].
$\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ is a structured sampler that exploits the first-order information of the potential (i.e., $\nabla\log(1/\pi)$) [girolami2011riemann]; its proposal is given as the solution to the Hamilton’s ODE equation, followed by the Metropolis-filter. In the linear-constraint setting above, this sampler requires $\mathcal{O}(md^{2/3}\log\frac{M}{\varepsilon})$ many iterations to achieve $\varepsilon$-close distance to $\pi^{\mathcal{K}}$ [lee2018convergence]. This sampler is further analyzed for practical ODE solvers [kook2022condition] and for more sophisticated self-concordant barriers [gatmiry2023sampling].
Similarly, $\mathsf{Mirror\ Langevin}$ [zhang2020wasserstein, jiang2021mirror, ahn2021efficient, li2022mirror] is a class of algorithms which converts the constrained problem into an unconstrained one obtained by considering the pushforward of the constrained space by $\nabla\phi$. The algorithm can also be metropolized [srinivasan2023fast]. The best known rates for this algorithm are $\widetilde{\mathcal{O}}(d\log\frac{1}{\varepsilon})$ under some strong assumptions on $\phi$.
@@
Samplers based on discretizations of Itô diffusions, stochastic processes which rapidly mix to $\pi$ in continuous time, have long been used for sampling without constraints [besag1995bayesian, dalalyan2012sparse, dalalyan2017further, chewi2023log]. While the underlying stochastic processes generalize easily to constrained settings, the discretization analysis relies crucially on the smoothness of the target distribution. This is clearly impossible to achieve in the constrained setting, and so some techniques are required to circumvent this difficulty. These algorithms, however, generalize easily to the more general problem of sampling from distributions of the form $\tilde{\pi}^{X}\propto e^{-f}\mathds{1}_{\mathcal{K}}$, by naturally incorporating first order information from $f$.
The first approach for adapting diffusion-based samplers [bubeck2015finite, bubeck2018sampling, lehec2023langevin] iterates a two-step procedure. First, a random step is taken, with $x_{k+1/2}\sim\mathcal{N}(x_{k},2hI_{d})$ for some appropriately chosen step $h$, [^4] A gradient step can be added in the more general case, for sampling from $\tilde{\pi}^{X}$ and then project it to $\mathcal{K}$, i.e., $x_{k+1}=\mathsf{proj}_{\mathcal{K}}(x_{k+1/2})$. The complexity is given in terms of queries to a projection oracle, each call to which can be implemented with a polynomial number of membership oracle queries; a total of $\tilde{\mathcal{O}}(\frac{d^{2}D^{3}}{\varepsilon^{4}})$ queries are needed to be $\varepsilon$-close in $\mathcal{W}_{2}$ to $\pi^{X}$. Another approach, which uses an algorithmically designed “soft” penalty instead of a projection, was proposed in [gurbuzbalaban2022penalized], and achieves a rate estimate of $\tilde{\mathcal{O}}(d/\varepsilon^{10})$.
A second approach, suggested by [pmlr-v65-brosse17a], considers a different proximal scheme, which performs a “soft projection” onto $\mathcal{K}$, by taking steps like $\mathcal{N}((1-h\lambda^{-1})x_{k}+h\mathsf{proj}_{\mathcal{K}}(x_{k}),2hI_{d})$. It is called Moreau-Yosida regularized Langevin, named after an analogous regularization scheme for constrained optimization. This scheme also relies on access to a projection oracle for $\mathcal{K}$, and quantifies their query complexity accordingly. Their final rate estimate is $\tilde{\mathcal{O}}(\frac{d^{5}}{\varepsilon^{6}})$ to be $\varepsilon$-close in $\mathsf{TV}$ distance to $\pi^{X}$.
@@
The Gibbs sampling scheme used in this paper was inspired by the restricted Gaussian oracle introduced in [lee2021structured] (in turn inspired by Gaussian Cooling [cousins2018gaussian]), which alternately iterates between a pure Gaussian step, and a “proximal” step (which we elaborate in our exposition). This scheme was given novel interpretations by [chen2022improved], which showed that it interpolates the forward and backward heat flows, in the sense defined by [klartag2021spectral]. The backward heat flow itself is intimately related to stochastic localization schemes, invented and popularized in [eldan2013thin, chen2021almost].
This formulation proved surprisingly powerful, allowing many existing rates in unconstrained sampling to be recovered from a relatively simple analysis. This was further extended by [fan2023improved] to achieve the current state-of-the-art rate in unconstrained sampling. Finally, [gopi2023algorithmic] suggest that this could be applied to tackle some constrained problems. However, the assumptions in this final mentioned work are not compatible with the uniform sampling problem on general convex bodies.

### v2

@@ 3.4 Putting it together
We can now show that $\mathsf{In\text{-}and\text{-}Out}$ subsumes previous results on uniform sampling from convex bodies (such as $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$), providing detailed versions of the main results in §1.2.
We first establish that the query complexity of $\mathsf{In\text{-}and\text{-}Out}$ matches that of the $\mathsf{Ball\ walk}$ under stronger divergences. Recall that $2\,\lVert\cdot\rVert_{\mathsf{TV}}^{2}\leq\mathsf{KL}\leq\log(1+\chi^{2})\leq\chi^{2}$.
@@
  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations. With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{\mathcal{O}}(qMd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total.
  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{\mathcal{O}}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{\mathcal{O}}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
@@
We just put together Lemma 14 and Theorem 23. For target accuracy $\varepsilon>0$, we use the $\mathcal{R}_{q}$-decay under (7) for $q\geq 2$ in Theorem 23. The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$. Due to $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 18), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$ iterations. Since each iteration has $\eta/m$-failure probability by Lemma 14, the union bound ensures that the total failure probability is at most $\eta$ throughout $m$ iterations. Lastly, each iteration requires $\widetilde{\mathcal{O}}(M\log^{4}\nicefrac{{1}}{{\eta\varepsilon}})$ membership queries in expectation by Lemma 14. Therefore, $\mathsf{In\text{-}and\text{-}Out}$ uses $\widetilde{\mathcal{O}}(qMd^{2}\min(D^{2},\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}})\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$ expected number of membership queries over $m$ iterations. Since $\mathcal{R}_{q}$ is non-decreasing in $q$, we can obtain the desired bound on $\mathcal{R}_{q}$ for $q\in[1,2)$.
For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\log d)$ (see Lemma 18). ∎
@@
$\mu_{1}^{X}(x)=\mathds{1}_{\mathcal{K}}(x)\cdot\int\frac{\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\exp(-\frac{1}{2h}\,\lVert y-x_{0}\rVert^{2})}{(2\pi h)^{d/2}\int_{\mathcal{K}}\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\,\mathrm{d}x}\,\mathrm{d}y\,.$
By Young’s inequality, $\lVert y-x\rVert^{2}\leq(\lVert y\rVert+D)^{2}\leq\frac{3}{2}\,\lVert y\rVert^{2}+3D^{2}$ for $x\in\mathcal{K}$. Hence,
$\int\frac{\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\exp(-\frac{1}{2h}\lVert y-x_{0}\rVert^{2})}{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\,\mathrm{d}x}\,\mathrm{d}y$
$\leq$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\,(\lVert y-x\rVert^{2}+\lVert y-x_{0}\rVert^{2}-\frac{3}{2}\lVert y\rVert^{2})\Bigr{)}\,\mathrm{d}y$
$=$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\bigl{(}\frac{1}{2}\lVert y-2(x+x_{0})\rVert^{2}+(\lVert x\rVert^{2}+\lVert x_{0}\rVert^{2}-2\lVert x+x_{0}\rVert^{2})\bigr{)}\Bigr{)}\,\mathrm{d}y$
$\leq$ $\frac{\exp(5h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{4h}\lVert y-2(x+x_{0})\rVert^{2}\Bigr{)}\,\mathrm{d}y$
@@
For $h=\tilde{\Theta}(d^{-2})$, $\mathsf{In\text{-}and\text{-}Out}$ requires $\widetilde{\mathcal{O}}(qd^{2}D^{2})$ iterations and in particular $\widetilde{\mathcal{O}}(qd^{2}D)$ iteration for isotropic uniform distributions. These results match those of $\mathsf{Speedy\ walk}$ [KLM06, LV17] (see Theorem 31).
@@ Acknowledgements.
We are deeply grateful to Andre Wibisono and Sinho Chewi for helpful comments and pointers to the literature for Lemma 19. This work was supported in part by NSF award 2106444, NSERC through the CGS-D award, and a Simons Investigator award.
@@ Appendix A Related work
Sampling from constrained log-concave distributions is a fundamental task arising in many fields. Uniform sampling with convex constraints is its simplest manifestation, which was first studied as a core subroutine for a randomized volume-computation algorithm [DFK91]. Since then, this fundamental problem has been studied for over three decades [LS90, LS93, KLS97, LV03, LV06, BEL18, BDMP17]. We review these algorithms, grouping them under three categories — geometric random walks, structured samplers, and diffusion-type samplers. Below, $\mathcal{K}$ is convex.
@@
We discuss two geometric random walks – $\mathsf{Ball\ walk}$ [LS93, KLS97] and $\mathsf{Hit\text{-}and\text{-}Run}$ [Smi84, Lov99]. $\mathsf{Ball\ walk}$ is a simple metropolized random walk; it draws $y$ uniformly at random from a ball of radius $\delta$ centered at a current point $x$, and moves to $y$ if $y\in\mathcal{K}$ and stays at $x$ otherwise. In the literature, $\mathsf{Ball\ walk}$ actually refers to a composite algorithm consisting of [$\mathsf{Speedy\ walk}+$ rejection sampling], where $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$ (see §B for details). The step size $\delta$ should be set to $\mathcal{O}(d^{-1/2})$ to avoid stepping outside of $\mathcal{K}$. [KLS97] showed that $\mathsf{Ball\ walk}$ needs $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ membership queries to be $\varepsilon$-close to $\pi^{\mathcal{K}}$ in $\mathsf{TV}$, where $D$ is the diameter of $\mathcal{K}$, and the warmness parameter $M$ measures the closeness of the initial distribution to the target uniform distribution $\pi^{\mathcal{K}}$.
$\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$. [LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lVert X-\mathbb{E}X\rVert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $\mathcal{O}(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries. Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to $\mathsf{Ball\ walk}$.
Both algorithm are affected by skewed shape of $\mathcal{K}$ (i.e., large $D$ or $R$), so these samplers are combined with pre-processing step called rounding. This procedure finds a linear transformation that makes the geometry of $\mathcal{K}$ less skewed and so more amenable to sampling. In literature, there exists a randomized algorithm [JLLV21] that rounds $\mathcal{K}$ and generates a good warm start (i.e., $M=\mathcal{O}(1)$), with $\mathsf{Ball\ walk}$ used as a core subroutine. We refer readers to [KZ24] for a streamlined proof, with $\mathsf{In\text{-}and\text{-}Out}$ used as a main sampler. This algorithm takes up $\widetilde{\mathcal{O}}(d^{3.5})$ queries in total, and in such position with the good warm start, $\mathsf{Ball\ walk}$ only needs $\widetilde{\mathcal{O}}(d^{2}\log\frac{1}{\varepsilon})$ queries to sample from $\pi^{\mathcal{K}}$.
@@
The first canonical example of such a zeroth-order sampler is $\mathsf{Dikin\ walk}$ used when $\mathcal{K}$ is given by $m$ linear constraints [KN12]; it draws a uniform sample from an ellipsoid (characterized by $\nabla^{2}\phi$) of fixed radius around a current point, and is often combined with a Metropolis adjustment. [KN12] shows that $\mathsf{Dikin\ walk}$ mixes in $\mathcal{O}(md\log\frac{M}{\varepsilon})$ steps, although each iteration is slightly more expensive than one membership query. This algorithm requires no rounding, but still needs a good warm-start, which can be achieved by an annealing-type algorithm using $\widetilde{\mathcal{O}}(md)$ iterations of $\mathsf{Dikin\ walk}$ [KV24].
$\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ is a structured sampler that exploits the first-order information of the potential (i.e., $\nabla\log(1/\pi)$) [GC11]; its proposal is given as the solution to the Hamilton’s ODE equation, followed by the Metropolis-filter. In the linear-constraint setting above, it requires $\mathcal{O}(md^{2/3}\log\frac{M}{\varepsilon})$ many iterations to achieve $\varepsilon$-close distance to $\pi^{\mathcal{K}}$ [LV18]. This sampler is further analyzed for practical ODE solvers [KLSV23] and for more sophisticated self-concordant barriers [GKV23].
Similarly, $\mathsf{Mirror\ Langevin}$ [ZPFP20, Jia21, AC21, LTVW22] is a class of algorithms which converts the constrained problem into an unconstrained one obtained by considering the pushforward of the constrained space by $\nabla\phi$. The algorithm can also be metropolized [SWW24]. The best known rate for this algorithm is $\widetilde{\mathcal{O}}(d\log\frac{1}{\varepsilon})$ under some strong assumptions on $\phi$.
@@
Samplers based on discretizations of Itô diffusions, stochastic processes which rapidly mix to $\pi$ in continuous time, have long been used for sampling without constraints [BGHM95, DT12, Dal17, Che24]. While the underlying stochastic processes generalize easily to constrained settings, the discretization analysis relies crucially on the smoothness of the target distribution. This is clearly impossible to achieve in the constrained setting, so some techniques are required to circumvent this difficulty. These algorithms, however, generalize easily to the more general problem of sampling from distributions of the form $\tilde{\pi}^{X}\propto e^{-f}\mathds{1}_{\mathcal{K}}$, by incorporating first-order information from $f$.
The first approach for adapting diffusion-based samplers [BEL15, BEL18, Leh23] iterates a two-step procedure. First, a random step is taken, with $x_{k+1/2}\sim\mathcal{N}(x_{k},2hI_{d})$ for some appropriately chosen step $h$, [^4] A gradient step can be added in the more general case, for sampling from $\tilde{\pi}^{X}$. and then project it to $\mathcal{K}$, i.e., $x_{k+1}=\mathsf{proj}_{\mathcal{K}}(x_{k+1/2})$. The complexity is given in terms of queries to a projection oracle, each call to which can be implemented with a polynomial number of membership oracle queries; a total of $\tilde{\mathcal{O}}(\nicefrac{{d^{2}D^{3}}}{{\varepsilon^{4}}})$ queries are needed to be $\varepsilon$-close in $\mathcal{W}_{2}$ to $\pi^{X}$. Another approach, which uses an algorithmically designed “soft” penalty instead of a projection, was proposed in [GHZ22], and achieves a rate estimate of $\tilde{\mathcal{O}}(\nicefrac{{d}}{{\varepsilon^{10}}})$.
A second approach, suggested by [BDMP17], considers a different proximal scheme, which performs a “soft projection” onto $\mathcal{K}$, by taking steps like $\mathcal{N}((1-h\lambda^{-1})x_{k}+h\mathsf{proj}_{\mathcal{K}}(x_{k}),2hI_{d})$. It is called Moreau-Yosida regularized Langevin, named after an analogous regularization scheme for constrained optimization. This scheme also relies on access to a projection oracle for $\mathcal{K}$, and quantifies their query complexity accordingly. Their final rate estimate is $\tilde{\mathcal{O}}(\nicefrac{{d^{5}}}{{\varepsilon^{6}}})$ to be $\varepsilon$-close in $\mathsf{TV}$ distance to $\pi^{X}$.
@@
The Gibbs sampling scheme used in this paper was inspired by the restricted Gaussian oracle introduced in [LST21] (in turn inspired by Gaussian Cooling [CV18]), which alternately iterates between a pure Gaussian step, and a “proximal” step (which we elaborate in our exposition). This scheme was given novel interpretations by [CCSW22], which showed that it interpolates the forward and backward heat flows, in the sense defined by [KP21]. The backward heat flow itself is intimately related to stochastic localization schemes, invented and popularized in [Eld13, Che21].
This formulation proved surprisingly powerful, allowing many existing rates in unconstrained sampling to be recovered from a relatively simple analysis. This was further extended by [FYC23] to achieve the current state-of-the-art rate in unconstrained sampling. Finally, [GLL+23] suggests that this could be applied to tackle some constrained problems. However, the assumptions in this final mentioned work are not compatible with the uniform sampling problem on general convex bodies.